A bookstore can purchase several calculators for a total cost of $200. If each calculator cost $1 less, the bookstore could purchase 10 additional calculators at the same total cost. How many calculators can be purchased at the regular price?
Answer by Boreal(15235)

the number is x
the cost is (200/x) for each
if the cost is $1 less, (200/x)-1, then they could buy 10 more or x+10, so the average cost of those would be (200/(x+10))
Therefore, (200/x)-1=200/(x+10)
or (200-x)/x=200/(x+10)
cross multiply
(200-x)(x+10)=200x
200x+2000-x^2-10x=200x
-x^2-10x+2000=0=x^2+10x-2000
this factors into (x-40)(x+50)=0
the only positive root is x=40 calculators ANSWER
each calculator is $5
if each calculator were $4, $1 less, they could buy 50 or 10 more.